2. Find the surface area of the object obtained s 2 about the y-axis. by rotating y: 478x2,15*
Answer #1

Given Sol and 1 < x <2 y= 4+372 Surface area of revolution about y-aris farads where ds = it + (dy je de Here, y= = 4+ 3x² dydet du 1+ 36 x²= u » d (1+36x²) = du 36(2) ada xda= 22 du And as x varies from 1 to 2 u varies from 37 to 145 2 145 és 277% V